### v2

@@ 2.1 Wagering Systems
2.2 Definitions
We start with an abstract definition of a prediction market. Our definition generalizes on recent computer science-based definitions [11, 30, 72] which each presume specific sub-types of a prediction market (e.g., [11, 72] use merging/splitting while [30] presumes automated bookmaking from §3.5). If our definitions are not clear, we refer the reader to Appendix 0.C where we describe a specific market offered by Polymarket and map each part of definitions to this real world example.
Definition 1(Market)
A (single) market is a tuple $M=(\Omega,J,R)$, where $\Omega$ is a nonempty outcome space, $J$ is a finite index set of contract labels (“shares”), and $R:\Omega\to\mathbb{R}_{\geq 0}^{J}$ is a nonnegative payoff map. Intuitively, $M$ is associated with an underlying uncertain event, whose possible resolutions are represented by $\Omega$. When $M$ resolves to $\omega\in\Omega$, one unit of share $j\in J$ pays $R(\omega)_{j}$ units of the numeraire $\mathcal{N}$ (Def. 2).
Remark: WTA special-case.
For a market $M=(\Omega,J,R)$, suppose there exists a bijection $\iota:J\to\Omega$ and $R(\omega)_{j}\in\{0,1\}$ with $\sum_{j\in J}R(\omega)_{j}=1$ for all $\omega\in\Omega$. Then $M$ is a winner-take-all (WTA) or Arrow–Debreu market: a unit claim of label $j$ pays $1$ iff the realized outcome equals $\iota(j)$, and 0 otherwise.
Definition 2(Prediction Market System)
A prediction–market system is a tuple $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathsf{res})$, where $\mathcal{M}$ is a countable set of markets, $\mathcal{N}$ is a numeraire (unit of account), and $\mathsf{res}$ is a resolution function such that, for each $M=(\Omega,J,R)\in\mathcal{M}$, we have $\mathsf{res}(M)\in\{\bot\}\cup\Omega$, $\mathsf{res}(M)$ is initially $\bot$, and $\mathsf{res}(M)$ transitions exactly once to some $\omega\in\Omega$.
Definition 3(System Axioms)
For every market $M=(\Omega,J,R)\in\mathcal{M}$ operating in system $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathsf{res})$, the following axioms hold. In what follows, all quantities are for this fixed market $M$, and we write $\mathsf{Treas}:=\mathsf{Treas}_{M}$ and $L:=L_{M}$.
  1. 1.
    Issuance and Solvency. Let $S=(S_{j})_{j\in J}\in\mathbb{R}_{\geq 0}^{J}$ be the outstanding supply vector, and let $\mathsf{Treas}\geq 0$ be the market’s treasury (in the numeraire $\mathcal{N}$). Define the worst-case liability
    $L(S)\;:=\;\sup_{\omega\in\Omega}\;\sum_{j\in J}S_{j}\,R(\omega)_{j}.$ (1)
    The system maintains, at all times,
    $\mathsf{Treas}\;\geq\;L(S).$ (2)
    Any operation that changes supplies (including issuance) must preserve solvency. In particular, if issuance changes supplies from $S$ to $S^{\prime}$, then treasury increases sufficiently:
    $\Delta\mathsf{Treas}\;\geq\;L(S^{\prime})-L(S).$ (3)
  2. 2.
    Transfer and Fungibility. Holdings are transferable between accounts and transfers conserve per–label totals $S_{j}$. Shares are fungible within each label: for all $j\in J$, any $\omega\in\Omega$, and any $q\geq 0$, redeeming $q$ units pays $q\,R(\omega)_{j}$ units of $\mathcal{N}$.
  3. 3.
    Settlement and redemption. If $\mathsf{res}(M)=\omega\in\Omega$, then any holder may redeem: for all $j\in J$ and $q\geq 0$, redeeming $q$ units of label $j$ pays $q\,R(\omega)_{j}$ units of $\mathcal{N}$. Redemption burns the redeemed shares and debits the treasury by the corresponding payout, preserving the solvency invariant above.
@@ 2.3 An Example of a Market
In game theory, cheap talk describes strategic misinformation or signalling aimed at shaping beliefs or prices, provided the cost of deception is outweighed by the potential payoff [21]. This is well illustrated by what followed in the HBO Satoshi market as new pieces of evidence emerged, some real and some fake, with some fakes relatively elaborate (professional appearing end-credits or hijacking a target’s X.com account) as summarized in Table 1. Further details are provided in Appendix 0.B.
@@ 3 Modular Workflow
Research Gap 2. Ca accuracy metrics [33] be used to compare various CePM and DePM platforms? Are there meaningful correlations between accuracy and the market microstructure options in Table 2?
@@ 3.3 Share Structure and Pricing
$R(\omega)_{j_{\mathsf{k}}}=\begin{cases}1,&\text{if }\omega=\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}$ (4)
For a WTA market to be well-functioning, two conditions must hold on outcome shares. (i) They should be mutually exclusive so no more than one share wins: $R(\omega)_{j_{\mathsf{k}}}\,R(\omega)_{j_{\mathsf{\ell}}}=0\quad\forall\omega\in\Omega,\forall\mathsf{k}\neq\mathsf{\ell}$; and (ii) they should be complete so at least one share wins: $\sum_{\mathsf{k}\in\Omega}R(\omega)_{j_{\mathsf{k}}}=1\quad\forall\omega\in\Omega$. If they are not mutually exclusive, the operator could be under-collateralized for making all payments (in violation of solvency in Definition 3). If they are incomplete, a deficient market might end with all participants receiving \$0. A consequence is that holding one share for each outcome is equivalent to holding \$1, a fact we will return to in §3.5.
@@
The second structure we term a yes-no bundle (YNB). YNB markets were popularized by InTrade [43]. A YNB market issues two share labels for each outcome, a ‘yes’ and a ‘no:’ $J=\{j_{\mathsf{A_{Y}}},j_{\mathsf{A_{N}}},j_{\mathsf{B_{Y}}},j_{\mathsf{B_{N}}},j_{\mathsf{C_{Y}}},j_{\mathsf{C_{N}}}\}$. For any $k\in\{\mathsf{A},\mathsf{B},\mathsf{C}\}$,
$R(\omega)_{j_{\mathsf{k_{Y}}}}=\begin{cases}1,&\text{if }\omega=\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}\qquad R(\omega)_{j_{\mathsf{k_{N}}}}=\begin{cases}1,&\text{if }\omega\neq\mathsf{k},\\ 0,&\text{otherwise.}\end{cases}$ (5)
Each outcome-specific pair $\{j_{\mathsf{k_{Y}}},j_{\mathsf{k_{N}}}\}$ constitutes a binary WTA market ($\mathsf{k}$ vs. not-$\mathsf{k}$). A YNB market is the union of these pairs, so the WTA exclusivity and completeness properties hold per pair. However exclusivity and completeness do not necessarily hold across all bundles, allowing more flexible markets. For example, a market on what words Trump will say in a congressional address included Bitcoin (no), beautiful at least 10 times (yes), and Canada (yes). [^15] Polymarket: ‘What will Trump say during address to Congress? Multiple words can resolve to yes (not exclusive) and it is possible he says none of the listed words (not complete). An established YNB market that is being actively traded, because it is not complete, can have new outcomes added to it fairly, whereas WTA markets must account for every possible outcome at the start of the market (perhaps utilizing an ‘other’ outcome).
@@
$\sum_{\mathsf{k}\in\Omega}R(\omega)_{j_{\mathsf{k_{Y}}}}=1\quad\text{and}\quad\sum_{\mathsf{k}\in\Omega}R(\omega)_{j_{\mathsf{k_{N}}}}=|\Omega|-1\quad\text{for all }\omega\in\Omega.$ (6)
The HBO documentary market from Section 2.3 is an example of a YNB-NR market (Polymarket markets are either YNB or YNB-NR [^16] Dune Analytics: ‘Polymarket - Activity and Volume’ compares YNB (called CTF there) and YNB-NR (called negrisk there) statistics.). This market was intended to be exclusive and complete across bundles by including a bundle for the outcome: ‘other/multiple.’ Holding a No outcome share for Hal Finney has the same payoff as holding a Yes share for every other candidate. Polymarket introduced a negRisk gadget that allows a trader to convert any No share into a portfolio of Yes shares for every other outcome. This enables traders to adjust their positions with less buying/selling on the markets, and also aligns prices between Yes and No markets with low friction arbitrage opportunities. Formally, a single No share has the equivalence (i.e., same payoff ignoring fees and discounting):
$R(\omega)_{j_{\mathsf{k_{N}}}}\;=\;\sum_{\ell\in\Omega\setminus\{\mathsf{k}\}}R(\omega)_{j_{\ell_{Y}}}\quad\text{for all }\omega\in\Omega\text{ and any }\mathsf{k}\in\Omega.$ (7)
While Yes has equivalence to:
$R(\omega)_{j_{k_{Y}}}\;=\;\sum_{\ell\in\Omega\setminus\{k\}}R(\omega)_{j_{\ell_{N}}}\;-\;(|\Omega|-2)$ (8)
@@ 3.4 Market Initialization
A near universal difference between any CePM and DePM is that a DePM allows outcome shares to be withdrawn from the platform, typically in a form compliant with a token standard such as ERC-20 or the more efficient ERC-1155 [52]. Withdrawing outcome shares allows traders to exchange tokens outside of the platform and to compose with third party DeFi services, including on-chain trading, lending, and leverage [82]. Options for trading outcome shares can be broken into two steps: (i) how does the first outcome share come into existence and how does the first trader trade, and (ii) how do traders trade once a market has been established? Probably the greatest evolution in DePMs, from Truthcoin to Polymarket, concerns how the first trade happens. There are three options: automated bookmaking, splitting, and matching.
@@
Fix a market $M=(\Omega,J,R)$. An automated bookmaker posts (i) a loss fund $\mathsf{Treas}\in\mathbb{R}_{\geq 0}$ in the numeraire $\mathcal{N}$ and (ii) a pricing rule (odds) for labels $j\in J$. It accepts deposits in $\mathcal{N}$ and issues shares, updating the supply vector $S$, and must always satisfy solvency (Definition 3).
@@
Fix a market $M=(\Omega,J,R)$ in numeraire $\mathcal{N}$. A splitting/merging mechanism provides counterparty-free state transitions for a fixed unit bundle $B$ of WTA share labels such that
@@
In YNB markets, each outcome’s YES/NO pair is its own two-outcome WTA market. Splitting is per outcome: converting \$1 of collateral mints one $j_{\mathsf{k_{Y}}}$ and one $j_{\mathsf{k_{N}}}$ for the chosen $\mathsf{k}$. Across outcomes, supplies are uncoupled—the total minted for the outcome $\mathsf{A}$’s bundle need not match that for outcomes $\mathsf{B}$ or $\mathsf{C}$.
@@ 3.8 Archiving
Research Gap 10. A useful tool that does not exist yet could use on-chain data, API data, and external data (news, social media, etc.) to replay a past market, allowing researchers to step forward and backward at each timestamp. It was onerous for us to manually study how how markets incorporate news (real and fake) into prices in §0.B without such a tool.
@@ Appendix 0.A List of DePMs
Appendix 0.B Satoshi HBO Market
Appendix 0.C Example instantiation of definition
In section 2.3, we discussed an example market concerning who the HBO documentary ‘Money Electric’ would name as Satoshi Nakamoto. In this section, we will see how this fits the definitions of a market, prediction market system, and the Arrow–Debreau special case. As discussed in Section 3.3, Polymarket employs a market mechanism we call a yes/no bundle (YNB), as opposed to winner-take-all (WTA). YNB requires an extra step in the definitions so we will do a first pass with a simplified WTA submarket, and then add the full YNB market.
0.C.1 Pass 1: Single WTA Market
Consider a simplified market that questions whether one specific candidate, e.g., Hal Finney, is named as Satoshi: yes or no. If through unforeseen circumstances, who the documentary names is not verifiable by the air date, the market resolves to no.
Recall Definition 1 of a market:
Definition 7(Market)
A (single) market is a tuple $M=(E,\Omega,J,R)$, where $E$ is a well-defined uncertain event, $\Omega$ is a nonempty outcome space for $E$, $J$ is a finite index set of contract labels (“shares”), and $R=(R_{j})_{j\in J}$ are nonnegative payoff functions with $R_{j}:\Omega\to\mathbb{R}_{\geq 0}$. We assume $|J|\geq|\Omega|$ and we require outcome distinguishability on $\Omega$:
$\forall\,\omega\neq\omega^{\prime}\in\Omega\ \ \exists\,j\in J:\ R_{j}(\omega)\neq R_{j}(\omega^{\prime}).$
When $M$ resolves to $\omega_{M}\in\Omega$, one unit of share $j\in J$ pays $R_{j}(\omega_{M})$ (in units of $\mathcal{N}$ defined below).
Event $E$ is whether or not Hal Finney is named as Satoshi in the documentary.
$\Omega$ is the set of resolution outcomes the market recognizes for $E$, the labels the system can publish at settlement. For this Hal-only binary market the outcome space is $\Omega=\{\mathsf{True},\mathsf{False}\}$. Here $\mathsf{True}$ means the documentary (per the market’s stated criteria) identifies Hal Finney as Satoshi; $\mathsf{False}$ aggregates all other possibilities (Hal not named, someone else named, no one named, the film does not air, or the identification is not verifiable by the resolution deadline).
We require that $\Omega$ contain no redundant labels. A label is redundant if it does not change at least one contract’s payoff: $\omega\sim\omega^{\prime}\iff\forall j\in J,\ R_{j}(\omega)=R_{j}(\omega^{\prime})$. For example, “Hal is named and it is raining” and “Hal is named and it is not raining” are distinct real-world states, but they cannot both appear in $\Omega$ since they both map to $\mathsf{True}$. The restriction can be written as:
$\forall\,\omega\neq\omega^{\prime}\in\Omega\ \ \exists\,j\in J:\ R_{j}(\omega)\neq R_{j}(\omega^{\prime}).$
In a prediction market, there are a set of shares. If we label them and add them all to an index set, that set is $J$. For this example, $J=\{\textsf{YES},\textsf{NO}\}$: Hal Finney is named (yes) and else (no). This is a normal case where each share in $J$ corresponds to an outcome in $\Omega$ but it is possible that the number of shares could exceed the number of outcomes. [^22] For example, consider a market of where Newcastle United (NUFC) finishes in the 2024-35 English Premier League season. Since there are 20 teams, the outcome has 20 possible labels: positions 1 to 20. Shares could exist for each of the 20 positions. But the outcome could also settle shares for whether NUFC finishes in the top 5 (which is relevant to champions league admittance), or shares on finishing in the bottom 3 (which is relevant to relegation).
$J$ is the index set of contract labels, the names of the tradeable shares. In this binary market we take $J=\{\textsf{YES},\textsf{NO}\}$.
The labels get their meaning from the component payoff functions $R_{j}:\Omega\to\mathbb{R}_{\geq 0}$. In this example, a payoff of 1 is given for shares that correctly predict the outcome and 0 otherwise. This means $R_{\textsf{YES}}(\mathsf{True})=1$, $R_{\textsf{YES}}(\mathsf{False})=0$, $R_{\textsf{NO}}(\mathsf{True})=0$, $R_{\textsf{NO}}(\mathsf{False})=1$.
Recall Definition 1 of a prediction–market system $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathsf{Res})$: $\mathcal{M}$ is the (countable) catalog of markets; $\mathcal{N}$ is the numeraire (unit of account used to price and settle claims); and $\mathsf{Res}=\{\mathrm{res}_{M}\}_{M\in\mathcal{M}}$ assigns to each market $M$ a resolution register that is initially $\bot$ and flips exactly once to some $\omega_{M}\in\Omega_{M}$. When $\mathrm{res}_{M}\neq\bot$, we set $\omega_{M}:=\mathrm{res}_{M}$ and each unit of label $j\in J$ settles for $R_{j}(\omega_{M})$ units of $\mathcal{N}$.
This market is a winner–take–all (Arrow–Debreu) special case: there is a bijection $\iota:J\to\Omega$ as follows: $\textsf{YES}\rightarrow\mathsf{True}$ and $\textsf{NO}\rightarrow\mathsf{False}$. Payoffs are $R_{j}(\omega)=\mathbf{1}\{\omega=\iota(j)\}$. Hence, for each $\omega\in\Omega$, exactly one label pays $1$ and all others pay 0.
The market tuple $M=(E,\Omega,J,R)$ specifies what to pay given an outcome (via $R$). The register $\mathrm{res}_{M}$ is the system’s single source of truth for which outcome actually occurred: before resolution $\mathrm{res}_{M}=\bot$ (no settlement), after resolution $\mathrm{res}_{M}=\omega_{M}\in\Omega_{M}$ (settlement applies).
Polymarket instantiates $\mathcal{S}$ with $\mathcal{M}$ equal to its live and historical markets, $\mathcal{N}$ the USD–denominated stablecoin USDC, and $\mathsf{Res}$ implemented by its on–chain resolution process (e.g., UMA’s optimistic oracle) that writes a single outcome to each $\mathrm{res}_{M}$.
For the Hal–only binary market, the system maintains a resolution register $\mathrm{res}_{M_{\mathrm{Hal}}}\in\{\bot\}\cup\Omega$ with $\Omega=\{\mathsf{True},\mathsf{False}\}$ (i.e., “yes/no” to the proposition). The register is initially $\bot$ and, after the platform’s resolution process completes, the oracle writes a single value $\omega_{M_{\mathrm{Hal}}}\in\Omega$ to the register.
Set $\omega_{M_{\mathrm{Hal}}}=\mathsf{True}$ iff the documentary (per stated criteria) identifies Hal Finney as Satoshi; otherwise set $\omega_{M_{\mathrm{Hal}}}=\mathsf{False}$.
Shares are fully collateralized to \$1 in the numeraire $\mathcal{N}$ (USDC): a unit of YES pays $1$ USDC at $\mathsf{True}$ and 0 USDC at $\mathsf{False}$; a unit of NO pays $1$ USDC at $\mathsf{False}$ and 0 USDC at $\mathsf{True}$. Formally,
$R_{\textsf{YES}}(\mathsf{True})=1,\ \ R_{\textsf{YES}}(\mathsf{False})=0,\qquad R_{\textsf{NO}}(\mathsf{True})=0,\ \ R_{\textsf{NO}}(\mathsf{False})=1.$
In the aired documentary, Peter Todd was named; therefore
$\mathrm{res}_{M_{\mathrm{Hal}}}=\omega_{M_{\mathrm{Hal}}}=\mathsf{False},$
and each unit settles as
$\textsf{YES}\ \to\ 0\ \text{USDC},\qquad\textsf{NO}\ \to\ 1\ \text{USDC}.$
0.C.2 Pass 2: YNB Market
We define a family of markets $\{M_{c}\}_{c\in C}$ where each $c\in C$ names one market in the family. For the HBO film, $C$ is the set of candidates including $\mathsf{Finney}$, $\mathsf{Szabo}$, $\mathsf{Sassaman}$, $\mathsf{Back}$, and $\mathsf{Other/Multiple}$. The event $E_{c}$ for a particular $\{M_{c}\}$ is: the documentary identifies $c$ as Satoshi. For each candidate’s market, the relevant market outcomes are $\Omega_{c}=\left\langle\mathsf{True},\mathsf{False}\right\rangle$. As a YNB market, Polymarket creates a yes share and a no share for each candidate $J_{c}=\left\langle\textsf{YES},\textsf{NO}\right\rangle$ called a yes/no bundle (YNB). The payoffs are as follows: $R^{(c)}_{\textsf{YES}}(\mathsf{True})=1$, $R^{(c)}_{\textsf{YES}}(\mathsf{False})=0$, $R^{(c)}_{\textsf{NO}}(\mathsf{True})=0$, and $R^{(c)}_{\textsf{NO}}(\mathsf{False})=1$. Each $\{M_{c}\}$ has its own volume of outstanding shares, its own pricing, and its own orderbook or AMM.
For settling, $\mathsf{Res}=\{\mathrm{res}_{M}\}_{M\in\mathcal{M}}$ gives each $M_{c}$ a register $\mathrm{res}_{M_{c}}\in\{\bot\}\cup\Omega_{c}$, initially $\bot$, that flips exactly once to $\omega_{M_{c}}\in\Omega_{c}$. The documentary named Peter Todd, who was not one of the named candidates and thus fell under Other/Multiple: $\omega_{M_{\textsf{Other/Multiple}}}=\mathsf{True}$ while $\omega_{M_{c}}=\mathsf{False}$ for all other $c\neq\textsf{Other/Multiple}$.
The $R^{(c)}_{j}(\omega_{M_{c}})$ for Other/Multiple: YES was $1$ USDC, Other/Multiple: NO was 0 USDC, every other named candidate’s NO paid $1$ USDC and their corresponding YES paid 0 USDC.

### v3

@@ 1 Introduction
Market Microstructure.
In traditional financial markets, the mechanics of trading are called market microstructure (distinguished from the study of prices). Our SoK covers design options for different stages of DePMs but the most depth is on microstructure topics, including how shares are structured, traded, and settled.
@@ 2.1 Wagering Systems
3 Example: Satoshi HBO Market
@@ 08 Oct Documentary airs and names Todd Very Significant Conclusive
In game theory, cheap talk describes strategic misinformation or signalling aimed at shaping beliefs or prices, provided the cost of deception is outweighed by the potential payoff [21]. This is well illustrated by what followed in the HBO Satoshi market as new pieces of evidence emerged, some real and some fake, with some fakes relatively elaborate (professional appearing end-credits or hijacking a target’s X.com account) as summarized in Table 1. Further details are provided in Appendix 3.
@@
3.1 Definitions
We start with an abstract definition of a DePM market, premised on outcome shares ($j\in J$) being represented as fungible token types with standard notions of accounts, transfers, minting (issuance), and burning (redemption). We do not try to capture CePMs designs that are margin-based (e.g., Kalshi [45]) where traders hold only account-level positions, although such systems often satisfy analogous solvency and settlement conditions at the level of accounts rather than transferable tokens. Our definition generalizes on recent computer science-based definitions [11, 30, 72] which each presume specific sub-types of a prediction market (e.g., [11, 72] use merging/splitting while [30] presumes automated bookmaking from §4.5).
Definition 1(Market)
A (single) market is a tuple $M=(\Omega,J,R,t_{\mathrm{res}})$, where $\Omega$ is a nonempty outcome space, $J$ is a finite index set of fungible token types, $R:\Omega\to\mathbb{R}_{\geq 0}^{J}$ is a nonnegative payoff map, and $t_{\mathrm{res}}\in\mathbb{N}$ is the (scheduled) resolution time. Intuitively, $M$ is associated with an underlying uncertain event, whose possible resolutions are represented by $\Omega$.
Remark: WTA special-case.
For a market $M=(\Omega,J,R,t_{\mathrm{res}})$, suppose there exists a bijection $\iota:J\to\Omega$ and $R(\omega)_{j}\in\{0,1\}$ with $\sum_{j\in J}R(\omega)_{j}=1$ for all $\omega\in\Omega$. Then $M$ is a winner-take-all (WTA) or Arrow-Debreu market: a unit claim of token $j$ pays $1$ iff the realized outcome equals $\iota(j)$, and 0 otherwise.
Definition 2(Prediction Market System)
A prediction market system is a tuple $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathrm{res})$, where $\mathcal{M}$ is a countable set of markets, $\mathcal{N}$ is a numeraire (unit of accounting), and $\mathrm{res}$ is a resolution function over discrete time $t\in\mathbb{N}$ such that, for each $M=(\Omega,J,R,t_{\mathrm{res}})\in\mathcal{M}$, we have $\mathrm{res}(M,t)\in\{\bot\}\cup\Omega$ for all $t\in\mathbb{N}$, $\mathrm{res}(M,t)=\bot$ for $t<t_{\mathrm{res}}$, and there exists $\omega\in\Omega$ such that $\mathrm{res}(M,t)=\omega$ for all $t\geq t_{\mathrm{res}}$.
Definition 3(System Axioms)
For every market $M=(\Omega,J,R,t_{\mathrm{res}})\in\mathcal{M}$ operating in system $\mathcal{S}=(\mathcal{M},\mathcal{N},\mathrm{res})$, the following axioms hold. In what follows, all quantities are for this fixed market $M$, and we write $\mathsf{Treas}:=\mathsf{Treas}_{M}$ and $L:=L_{M}$.
  1. 1.
    Issuance and Solvency.
    Let $Q=(Q_{j})_{j\in J}\in\mathbb{R}_{\geq 0}^{J}$ be the vector of outstanding supplies, where $Q_{j}$ is the total number of units of token $j$ that have been issued and not yet burned, and let $\mathsf{Treas}\geq 0$ be the market’s treasury (in the numeraire $\mathcal{N}$).
    Define the worst-case liability
    $\mathrm{L}(Q)\;:=\;\sup_{\omega\in\Omega}\;\sum_{j\in J}Q_{j}\,R(\omega)_{j}.$ (1)
    The system maintains, at all times,
    $\mathsf{Treas}\;\geq\;\mathrm{L}(Q).$ (2)
    Any operation that changes supplies (including issuance) must preserve solvency. In particular, if issuance changes supplies from $Q$ to $Q^{\prime}$, then the treasury is adjusted sufficiently:
    $\Delta\mathsf{Treas}\;\geq\;\mathrm{L}(Q^{\prime})-\mathrm{L}(Q).$ (3)
  2. 2.
    Transfer and Fungibility. Holdings are transferable between accounts and transfers conserve per-token totals $Q_{j}$ (the system-wide outstanding supply of token $j$).
    Units of each token type are fungible: for all $j\in J$, any $\omega\in\Omega$, and any $\geq 0, aa$ units of token $j$ entitle the holder to $a\,R(\omega)_{j}$ units of $\mathcal{N}$ if the realized outcome is $\omega$.
  3. 3.
    Settlement and Redemption. If $t\geq t_{\mathrm{res}}$ and $\mathrm{res}(M,t)=\omega\in\Omega$, then any holder may redeem: for all $j\in J$ and $a\geq 0$, redeeming $a$ units of token $j$ pays $a\,R(\omega)_{j}$ units of $\mathcal{N}$. Redemption burns the redeemed tokens and debits the treasury by the corresponding payout, preserving the solvency invariant above.
@@ 4 Modular Workflow
Research Gap 2. Can accuracy metrics [33] be used to compare various CePM and DePM platforms? Are there meaningful correlations between accuracy and the market microstructure options in Table 2?
@@ 4.1 Underlying Infrastructure
A near universal difference between any CePM and DePM is that a DePM allows outcome shares to be withdrawn from the platform, typically in a form compliant with a token standard such as ERC-20 or the more efficient ERC-1155 [52]. Withdrawing outcome shares allows traders to exchange tokens outside of the platform and to compose with third party DeFi services, including on-chain trading, lending, and leverage [82].
@@ 4.3 Share Structure and Pricing
$R(\omega)_{j_{k}}=\begin{cases}1,&\text{if }\omega=k,\\ 0,&\text{otherwise.}\end{cases}$ (4)
For a WTA market to be well-functioning, two conditions must hold on outcome shares. (i) They should be mutually exclusive so no more than one share wins: $R(\omega)_{j_{k}}\,R(\omega)_{j_{\mathsf{\ell}}}=0\quad\forall\omega\in\Omega,\forall k\neq\mathsf{\ell}$; and (ii) they should be complete so at least one share wins: $\sum_{k\in\Omega}R(\omega)_{j_{k}}=1\quad\forall\omega\in\Omega$. If they are not mutually exclusive, the operator could be under-collateralized for making all payments (in violation of solvency in Definition 3). If they are incomplete, a deficient market might end with all participants receiving \$0. A consequence is that holding one share for each outcome is equivalent to holding \$1, a fact we will return to in §4.5.
@@
The second structure we term a yes-no bundle (YNB). YNB markets were popularized by InTrade [43]. A YNB market issues two token types for each outcome, a ‘yes’ and a ‘no:’ $J=\{j_{\mathsf{A_{Y}}},j_{\mathsf{A_{N}}},j_{\mathsf{B_{Y}}},j_{\mathsf{B_{N}}},j_{\mathsf{C_{Y}}},j_{\mathsf{C_{N}}}\}$. For any $k\in\{\mathsf{A},\mathsf{B},\mathsf{C}\}$,
$R(\omega)_{j_{\mathsf{k_{Y}}}}=\begin{cases}1,&\text{if }\omega=k,\\ 0,&\text{otherwise.}\end{cases}\qquad R(\omega)_{j_{\mathsf{k_{N}}}}=\begin{cases}1,&\text{if }\omega\neq k,\\ 0,&\text{otherwise.}\end{cases}$ (5)
Each outcome-specific pair $\{j_{\mathsf{k_{Y}}},j_{\mathsf{k_{N}}}\}$ constitutes a binary WTA market ($k$ vs. not-$k$). A YNB market is the union of these pairs, so the WTA exclusivity and completeness properties hold per pair. However exclusivity and completeness do not necessarily hold across all bundles, allowing more flexible markets. For example, a market on what words Trump will say in a congressional address included Bitcoin (no), beautiful at least 10 times (yes), and Canada (yes). [^15] Polymarket: ‘What will Trump say during address to Congress? Multiple words can resolve to yes (not exclusive) and it is possible he says none of the listed words (not complete). An established YNB market that is being actively traded, because it is not complete, can have new outcomes added to it fairly, whereas WTA markets must account for every possible outcome at the start of the market (perhaps utilizing an ‘other’ outcome).
@@
$\sum_{k\in\Omega}R(\omega)_{j_{\mathsf{k_{Y}}}}=1\quad\text{and}\quad\sum_{k\in\Omega}R(\omega)_{j_{\mathsf{k_{N}}}}=|\Omega|-1\quad\text{for all }\omega\in\Omega.$ (6)
The HBO documentary market from Section 3 is an example of a YNB-NR market (Polymarket markets are either YNB or YNB-NR [^16] Dune Analytics: ‘Polymarket - Activity and Volume’ compares YNB (called CTF there) and YNB-NR (called negrisk there) statistics.). This market was intended to be exclusive and complete across bundles by including a bundle for the outcome: ‘other/multiple.’ Holding a No outcome share for Hal Finney has the same payoff as holding a Yes share for every other candidate. Polymarket introduced a negRisk gadget that allows a trader to convert any No share into a portfolio of Yes shares for every other outcome. This enables traders to adjust their positions with less buying/selling on the markets, and aligns prices between Yes and No markets with low friction arbitrage opportunities. Formally, a single No share has the equivalence (i.e., same payoff ignoring fees and discounting):
$R(\omega)_{j_{\mathsf{k_{N}}}}\;=\;\sum_{\ell\in\Omega\setminus\{k\}}R(\omega)_{j_{\ell_{Y}}}\quad\text{for all }\omega\in\Omega\text{ and any }k\in\Omega.$ (7)
Yes has equivalence to:
$R(\omega)_{j_{\mathsf{k_{Y}}}}=\sum_{\ell\in\Omega\setminus\{k\}}R(\omega)_{j_{\ell_{N}}}-(|\Omega|-2)$ (8)
NegRisk also allows a trader holding a portfolio of No shares across $A$ outcomes ($|A|>0$) to extract cash by converting to the complimentary Yes shares.
$\sum_{k\in A}R(\omega)_{j_{k_{N}}}=(|A|-1)+\sum_{\ell\in\Omega\setminus A}R(\omega)_{j_{\ell_{Y}}}\quad\text{for all }\omega\in\Omega\text{ and any }A\subseteq\Omega.$ (9)
Example.
Alice owns 1000 No shares for Adam Back, Hal Finney, and Nick Szabo in the HBO documentary set to identify Satoshi Nakamoto. Since only one share type can win, Alice is guaranteed at least \$2000 from her positions at resolution (and up to \$3000 is she is right about all three). It is capital inefficient to keep her \$2000 locked up in her portfolio of No shares. NegRisk allows her to converts her position into 1000 Yes shares for every other candidate and extract the \$2000 in cash, maintaining the exact same exposure to the market (\$1000 at risk).
@@ 4.4 Market Initialization
Options for trading outcome shares can be broken into two steps: (i) how does the first outcome share come into existence and how does the first trader trade, and (ii) how do traders trade once a market has been established? Probably the greatest evolution in DePMs, from Truthcoin to Polymarket, concerns how the first trade happens. There are three options: automated bookmaking, splitting, and matching.
@@
Fix a market $M=(\Omega,J,R,t_{\mathrm{res}})$. An automated bookmaker posts (i) a loss fund $\mathsf{Treas}\in\mathbb{R}_{\geq 0}$ in the numeraire $\mathcal{N}$ and (ii) a pricing rule (odds) for token types $j\in J$. It accepts deposits in $\mathcal{N}$ and issues shares, updating the supply vector $S$, and must always satisfy solvency (Definition 3).
@@
Fix a market $M=(\Omega,J,R,t_{\mathrm{res}})$ in numeraire $\mathcal{N}$. A splitting/merging mechanism provides counterparty-free state transitions for a fixed unit bundle $B$ of WTA token types such that
@@
In YNB markets, each outcome’s YES/NO pair is its own two-outcome WTA market. Splitting is per outcome: converting \$1 of collateral mints one $j_{\mathsf{k_{Y}}}$ and one $j_{\mathsf{k_{N}}}$ for the chosen $k$. Across outcomes, supplies are uncoupled—the total minted for the outcome $\mathsf{A}$’s bundle need not match that for outcomes $\mathsf{B}$ or $\mathsf{C}$.
@@ 4.8 Archiving
Research Gap 10. A useful tool that does not exist yet could use on-chain data, API data, and external data (news, social media, etc.) to replay a past market, allowing researchers to step forward and backward at each timestamp. It was onerous for us to manually study how how markets incorporate news (real and fake) into prices in §3 without such a tool. It would also make it possible to backtest and train trading strategies using strictly point-in-time information, avoiding hindsight bias from revised data or retrospectively curated search results.
Research Gap 11. A useful feature that does not exist yet would add an automated feedback loop after markets resolve. The system could analyze what happened (using on-chain data, platform logs, and external signals) and use it to tune future markets, such as updating fees, incentives, liquidity settings, or oracle parameters. Today these updates are mostly manual or governance-driven.